OpenStudy (anonymous):

what is the inverse of this function X^2 +5x-6 if x<0 I am struggling with the x^2 + 5x part please help

OpenStudy (amorfide):

let f(x)= X^2 +5x-6 let y=f(x) y= X^2 +5x-6 complete the square and make x the subject

OpenStudy (anonymous):

I don't know how to complete the square the rest I can do the the spuare is confusing

OpenStudy (anonymous):

i know x^2+5x-6 = (x+6) (x-1)

OpenStudy (amorfide):

you factorised it, which will not help us, if we complete the square we can get x on its own. to complete the square you want to do x plus half the coefficient of x square it then take away the square so like this... X^2 +5x-6 you do (x+5/2)^2 - (5/2)^2 -6 the negative 6 just stays there so now you have y=(x+5/2)^2 -(5/2)^2 - 6 can you do this from here?

OpenStudy (anonymous):

then the last step is (x+2.5)^2-12.25 ?

OpenStudy (amorfide):

yes then make x the subject once you do that, you replace the x with f^1(x) and replace your y with x to get your inverse function I must go to school now, but will check what answer you get in 30 mins

OpenStudy (anonymous):

thanks a lot x= (9y+12.25)^0.5 -2.5

OpenStudy (amorfide):

why did you get 9y?

OpenStudy (amorfide):

you should get x=(y+12.25)^0.5 -2.5 then finish of by replacing x with f^-1(x) and replacing y with x to get F^-1(x)= (x+12.25)^0.5 -2.5

OpenStudy (anonymous):

thanks for all the trouble